OpenStudy (anonymous):

The demand for concert tickets depends on the price per ticket. Thus, the demand is a function of the price. If this function is called D(p), use the formula below to compute D(15), the demand for tickets when the price is $15. D(p)=2115-2p-p^2 A. D(15)=1120 B. D(15)=2115-225 C. D(15)=2115-60 D. D(15)=2115-255 (Please help)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

D(p) = 2115 - 2p - p^2 D(15) = 2115 - 2(15) - 15^2 {substituted 15 in for p} D(15) = 2115 - 30 - 225 D(15) = 2115 - 255 {combined -30 and -225} Answer is: C

OpenStudy (anonymous):

so C or D?